40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 799     Entry time: Tue Aug 5 12:52:28 2008
Author: Yoichi 
Type: Update 
Category: SUS 
Subject: ITMX, SRM OSEM spectra 
Free swinging spectra of ITMX and SRM.
ITMX seems to be ok after yesterday's work, though the OSEM DC values are still a bit off from the normal value of 0.9.
(ITMX OSEM values: UL=1.12, UR=1.38, LR=0.66, LL=0.41, SIDE=0.66)
SRM is still clearly wrong.
Attachment 1: ITMX-2008_08_05-morning.pdf  162 kB  | Hide | Hide all
ITMX-2008_08_05-morning.pdf
Attachment 2: SRM-2008_08_05-morning.pdf  180 kB  | Hide | Hide all
SRM-2008_08_05-morning.pdf
ELOG V3.1.3-